1. What is the parent nuclide that is bombarded with an alpha particle to generate curium-242 (Cm-242) and a neutron?

Cf-249

Md-256

Pu-239

Es-253

U-238
Answer #1

23 Pu + He +96 242 Cm +on

From above reaction , it is clear that the parent nucleus is Pu-239

total mass number on reactant side must be = total mass number on product side

239 +4 = 243

242+1 = 243

total atomic number on reactant side = total atomic number on product side

94+2 =94

94+0 =94

thus answer

Pu-239